# A Review of Traffic Wave Suppression Strategies: Variable Speed Limit vs. Jam-Absorption Driving

### v1

@@ III-B Traffic Dynamics Modeling
Recently, the rapidly-emerged Generative AI appears to be a potential tool to mitigate these challenges. An insightful example might be PromptGAT [74], which integrates Large Language Models (LLMs) to address the Sim-to-Real transfer performance gap in traffic signal control. The model utilizes the reasoning capabilities of LLMs to generate dynamic information based on specific conditions, such as weather, traffic status, and road types, which are then applied to enhance the action transformation process in RL models. This approach improves the applicability of simulation-based strategies in real-world environments, suggesting that LLMs can help theoretical models better adapt to real-world settings, which are often complex and challenging for humans.
Additionally, in some real-time applications, accurately determining or predicting which assumption best matches the actual situation is still challenging. This is not only due to the dynamic and stochastic nature of real-world conditions, but also because some information, particularly global information, cannot be obtained solely through observing local data. As reported in [21], the proposed VSL strategy was activated in approximately half of the cases for stop-and-go waves, and in the other half for different types of jams. It unveils the difficulty of identifying qualified stop-and-go waves in practice, i.e., distinguishing an isolated wave from multiple waves with certain criteria. Similarly, to quickly determine if a stop-and-go wave has consistently formed, [54] had to propose a two-step estimation approach, which allows for an adjustment of the estimation after the initial judgment.
III-C Traffic State Estimation and Prediction
As pointed out by most of the current research, traffic state estimation and prediction are the foundation of successful execution of wave suppression. Despite the importance, most of the current works do not incorporate state-of-the-art approaches. For instance, the current MPC-based VSL strategies heavily rely on the prediction of cell-based models such as METANET in [39, 40] and CTM in [41, 70]. In addition to the inherent limitations of model-based approaches (such as oversimplification), the resolution requirements of those cell-based models seem to be overlooked. Specifically, the hard constraint for those models that guarantees conservation laws (i.e., all flow falls in a cell at least once) is $L_{\text{C}}>v_{\text{f}}T_{\text{C}}$, where $v_{\text{f}}$ is the free-flow speed, and $L_{\text{C}}$ and $T_{\text{C}}$ are the length (in space) and width (in time) of a cell, respectively (Figure 9(a)). The typical free-flow speed on a freeway is 120 km/h. Therefore, for a cell with a 20-second time width, the corresponding length must exceed 600 meters, which may be too coarse to effectively capture traffic dynamics in practice, particularly in multiple stop-and-go waves [12].
@@ III-C Traffic State Estimation and Prediction
In addition, it is worth noting that, beyond accurate prediction of stop-and-go wave propagation, inflow or demand prediction is also critical for effective suppression, as reported by [80]. However, this aspect appears to be largely overlooked in most existing studies, which often assume solely a constant and typically high inflow that sustains stop-and-go waves.
III-D Stochasticity
Most of the theory-based strategies or analytical models rely on a theory that can describe traffic dynamics, such as KWT at the macroscopic level and CF models at the microscopic level. Although these fundamental theories play an important role in modeling and analysis, the stochastic nature of traffic indicates that these oversimplified deterministic models, which basically describe an equilibrium or average behavior, may not be sufficient to support practical applications. Therefore, it is necessary to adapt existing strategies to encompass stochastic extension of the current fundamental theories. For instance, [81] extended KWT by incorporating stochastic FDs and stochastic initial conditions, which can be employed to enhance the current KWT-based strategy and analysis, such as stochastic versions of SPECIALIST [44] and the theory of JAD [51, 56]. Similarly, stochastic CF models, such as those in [82, 83], can be used to derive the conditions for stop-and-go wave suppression and the prevention of secondary waves [52, 53]. Among the existing VSL and JAD works, [57] provides a valuable example, as it attempts to randomize a deterministic CF model and incorporate it into the existing JAD framework for stop-and-go wave propagation estimation. A similar technique could also be applied in other aspects of the overall framework of stop-and-go wave suppression.
@@ III-D Stochasticity
As a rapidly developing field, RL deserves more exploration. To date, only one RL-based strategy has been identified, and it is only for VSL [20]. There is significant potential for follow-up research, such as adopting a microscopic perspective or introducing RL into JAD, as [57] introduced MPC into JAD. In this context, it would be valuable to consider stochastic factors, such as heterogeneous driving behavior and human drivers’ imperfection [8].
In testing a strategy, although simulation is unavoidable most of the time, the more realistic the simulation is, the more convincing the strategy will be and the closer it will be to practical deployment. Those stochastic (particularly microscopic) models, such as those advanced models mentioned in [86, 87], are expected to be employed, while most of the current strategies were tested in a deterministic or aggregated model such as IDM, METANET, and CTM. It is worth mentioning that many data-driven traffic flow models have recently emerged [86, 87]. Since these models are independent of any specific mathematical framework and incorporate randomness, using them to test a strategy can lead to a promising validation [54, 57].
III-E Scenarios for Strategy Validation
Figure 10: JAD performance in a more challenging multiple-wave scenario (sourced from [54]). Both (a) and (b) share the same inflow. Fewer waves are triggered in (a), while more in (b).
As shown in Figure 1, it is quite common to observe that multiple and periodic stop-and-go waves appear on the freeway at the same time, and therefore, it is necessary to evaluate the strategies in a more practical and common scenario. Unfortunately, as shown in Table III, many strategies are validated in single-wave scenarios, which are relatively simple compared to the real-world observations. Single-wave scenarios are well-suited for initially testing the strategies, while multiple-wave scenarios allow for a more thorough assessment of the strategies in a setting closer to real-world conditions. Obviously, multiple-wave scenarios present greater challenges (Figure 10) [54, 56]. For the cell-based MPC strategies, the model resolution, i.e., the cell size, may not be sufficiently fine to distinguish multiple waves. However, it remains uncertain whether the current cell size is sufficient to capture multiple waves, or what level of granularity would be appropriate. Even for the SPECIALIST, although it was designed for the single (or isolated) wave scenario, it still deserves a test in a more realistic scenario such as the one shown in Figure 1(d). Then, new problems can be uncovered, inspiring new research questions and corresponding solutions.
@@ III-E Scenarios for Strategy Validation
III-F Field Test and Practical Deployment
III-F1 Variable Speed Limit
@@
Figure 11: Two ways to disseminate speed limits (The background image is sourced from [21]): (a) One-shot plan generated by SPECIALIST, where the VSL display across multiple VMS is scheduled in advance based on the time-space plan shown in the figure; (b) Two-VMS configuration.
At present, the VSL strategies require the installation of a series of gantries and VMS for support. Referring to JAD strategies, it may be feasible to propose a low-cost bi-VSL strategy, in which the upstream VSL reduces the speed limit, while the downstream one is responsible for setting the speed back to normal (Figure 11(b)). The bi-location VSL might not only be more practical but also sufficient to set a low-speed region to prevent stop-and-go wave propagation. One may not need to worry too much about the compliance rate, as even if only a small portion of drivers respond to the lower speed limit, the overall traffic will slow down.
III-F2 Jam Absorption Driving
@@
First, the behavior similar to JAD has occurred in the real world. As shown in Figure 12, a police car was recorded swerving on the freeway in California, USA, to control the following traffic at a high speed. In Figure 12, it was 70 km/h, which is higher than the lowest VSL at 60 km/h required by many VSL studies. The real-world police car swerving highlights the feasibility of JAD. Although the goal of the police car in this example might not be jam absorption, it greatly inspires a potential real-world implementation of JAD.
@@
Second, empirical data suggest that maintaining an appropriate gap from the leading vehicle when joining a stop-and-go wave could be effective in suppressing it, as shown in Figure 13. Although one may argue that this observation is related to the drop in demand, it still implies the potential feasibility of JAD, in particular from the microscopic perspective. The key lies in identifying the applicable conditions under which JAD can be effectively implemented. Therefore, it suggests another promising research direction, which is to extract similar scenarios from the existing extensive empirical trajectory datasets and analyze them to identify the conditions under which such behavior could contribute to wave suppression, ultimately proposing an empirical observation-based JAD strategy or even other CAV-based strategies beyond JAD.
Figure 13: Empirical observations suggest that maintaining an appropriate gap from the leading vehicle when joining a stop-and-go wave may effectively suppress its propagation. The regional screenshots were extracted from time-space diagrams of vehicle trajectories, which were plotted using data collected from the Hanshin Expressway Route 11 (Ikeda Route) in Osaka City, Japan [18].
@@ IV Concluding Remarks
Stop-and-go waves are the common pattern of congestion on freeways/highways worldwide. With their severe adverse effects on traffic efficiency, driving safety, and emissions, stop-and-go waves persistently affect human society on a broad time-space scale. Over decades of research and development, a variety of measures and strategies have emerged to suppress the propagation of stop-and-go waves, including mainly VSL and JAD strategies. Unfortunately, despite important similarities, such as the shared goals and underlying logic, a disconnection between the two areas is observed. On the occasion of the 20th anniversary of the first stop-and-go wave suppression strategy [39] being proposed, this review summarizes the progress of both VSL and JAD areas in detail, pinpoints the strengths and weaknesses of each, and seeks to bridge the two areas, identify research opportunities, and streamline research directions.
Learning from JAD strategies, we recommend that the VSL area to place more focus on thorough validation using more realistic (e.g., stochastic) microscopic traffic simulations and in more realistic scenarios. The secondary wave issue, which has been widely reported in the JAD area and is highly likely to exist in the VSL area, urgently needs to be tested and addressed. To enhance traffic estimation and prediction, more advanced techniques, such as deep learning, could be incorporated. To increase the generalization of the strategies, it is beneficial to consider more diverse traffic patterns, as well as introduce the latest AI tools (e.g., LLMs) to reduce the workload of fine-tuning. To reduce the need for constructing continuous, high-cost gantries for VSL strategies, a lightweight bi-VSL strategy could be proposed, drawing inspiration from JAD strategies.
Compared to VSL strategies, the lack of solid theoretical foundations in JAD strategies is more prominently exposed. This may be the main reason why the secondary wave issue has been inadequately addressed, and some research may even be biased. The real-world test in [21] provides strong evidence for the effectiveness of VSL strategies. Unfortunately, no similar field test for JAD has been reported. Since they share similar logic, JAD should also work similarly to VSL. Although it is not directly related to stop-and-go wave suppression, the observation of a police car swerving on a freeway at high speeds provides us with great confidence that JAD is also practically feasible. Making the current requirements (i.e., model input) for JAD implementation more practical is a valuable research effort that we can make before a real-world test opportunity arises.
@@
Unlike other areas of transportation research that center on understanding mechanisms or identifying patterns, freeway traffic control directly interacts with the traffic system, placing exceptionally high demands on the robustness and adaptability of the models and strategies. Given the similarities in research questions and goals, it might be worth exploring a unified framework that decouples modeling from implementation. Specifically, at the theoretical level, we may focus on stop-and-go wave suppression-oriented traffic flow optimization, through, e.g., traffic-based optimal control with the goal of minimizing travel time or suppressing the waves. In implementation, choosing between VSL and JAD can be guided by their respective strengths and limitations, depending on the practical conditions, with the aim of achieving the most effective outcome. Such a plug-in modeling-implementation structure enhances the overall framework’s generality and flexibility.

### v2

@@ III-C Generalizability of Strategies
Recently, the rapidly-emerged Generative AI appears to be a potential tool to mitigate these challenges. An insightful example might be PromptGAT [74], which integrates Large Language Models (LLMs) to address the Sim-to-Real transfer performance gap in traffic signal control at intersections. The model utilizes the reasoning capabilities of LLMs to generate dynamic information based on specific conditions, such as weather, traffic status, and road types, which are then applied to enhance the action transformation process in RL models. This approach improves the applicability of simulation-based strategies in real-world environments, suggesting that LLMs can help theoretical models better adapt to real-world settings, which are often complex and challenging for humans.
Additionally, in some real-time applications, accurately determining or predicting which assumption best matches the actual situation is still challenging. This is not only due to the dynamic and stochastic nature of real-world conditions, but also because some information, particularly global information, cannot be obtained solely through observing local data. As reported in [21], the proposed VSL strategy was activated in approximately half of the cases for isolated stop-and-go waves, and in the other half for different types of jams. It unveils the difficulty of identifying qualified stop-and-go waves in practice, i.e., distinguishing an isolated wave from multiple waves with certain criteria. Similarly, to quickly determine if a stop-and-go wave has consistently formed, [54] had to propose a two-step estimation approach, which allows for an adjustment of the estimation after the initial judgment.
III-D Traffic State Estimation and Prediction
As pointed out by most of the current research, traffic state estimation and prediction are the foundation of the successful execution of wave suppression. Despite the importance, most of the current works do not incorporate state-of-the-art approaches. For instance, the current MPC-based VSL strategies heavily rely on the prediction of cell-based models such as METANET in [39, 40] and CTM in [41, 70]. In addition to the inherent limitations of model-based approaches (such as oversimplification), the resolution requirements of those cell-based models seem to be overlooked. Specifically, the hard constraint for those models that guarantees conservation laws (i.e., all flow falls in a cell at least once) is $L_{\text{C}}>v_{\text{f}}T_{\text{C}}$, where $v_{\text{f}}$ is the free-flow speed, and $L_{\text{C}}$ and $T_{\text{C}}$ are the length (in space) and width (in time) of a cell, respectively (Figure 9(a)). The typical free-flow speed on a freeway is 120 km/h. Therefore, for a cell with a 10-second time width (typically used by METANET), the corresponding length must exceed 300 meters, which may be too coarse to effectively capture traffic dynamics with the wave speed of around 16 km/h in practice, particularly in multiple stop-and-go waves [12].
@@ III-D Traffic State Estimation and Prediction
It is worth noting that, beyond accurate prediction of stop-and-go wave propagation, inflow or demand prediction is also critical for effective suppression, as reported by [80]. However, this aspect appears to be largely overlooked in most existing studies, which often assume solely a constant and typically high inflow that sustains stop-and-go waves.
III-E Robustness to Randomness
Most of the theory-based strategies or analytical models rely on a theory that can describe traffic dynamics, such as KWT at the macroscopic level and CF models at the microscopic level. Although these fundamental theories play an important role in modeling and analysis, the stochastic nature of traffic indicates that these oversimplified deterministic models, which basically describe an equilibrium or average behavior, may not be sufficient to support practical applications. Therefore, it is necessary to adapt existing strategies to encompass stochastic extension of the current fundamental theories. For instance, [81] extended KWT by incorporating stochastic FDs and stochastic initial conditions, which can be employed to enhance the current KWT-based strategy and analysis, such as considering stochastic factors when tuning SPECIALIST [44] and enhancing the theory of JAD [51, 56]. Similarly, stochastic CF models, such as those in [82, 83], can be used to derive the conditions for stop-and-go wave suppression and the prevention of secondary waves [52, 53]. Among the existing VSL and JAD works, [57] provides a valuable example, as it attempts to randomize a deterministic CF model and incorporate it into the existing JAD framework for wave propagation estimation. A similar technique could also be applied in other aspects of the overall framework of stop-and-go wave suppression.
@@ III-E Robustness to Randomness
As a rapidly developing field, RL warrants further exploration for wave suppression tasks, given its inherent capability to handle stochasticity and non-stationary dynamics. To date, only one RL-based strategy has been identified, and it is only for VSL [20]. There is significant potential for follow-up research, such as adopting a microscopic perspective or introducing RL into JAD, as [57] introduced MPC into JAD. In this context, it would be valuable to consider stochastic factors, such as heterogeneous driving behavior and human drivers’ imperfection [8].
In testing a strategy, although simulation is unavoidable most of the time, the more realistic the simulation is, the more convincing the strategy will be and the closer it will be to practical deployment. In particular, accurately reproducing the capacity drop—preferably aligned with the findings of [86] on the relationship between queue discharge rate and in-queue speed—as well as the instability and breakdown phenomena at bottlenecks, is essential. Those stochastic (particularly microscopic) models, such as those advanced models mentioned in [87, 88], are expected to be employed, while most of the current strategies were tested in a deterministic or aggregated model such as IDM, METANET, and CTM. It is worth mentioning that many data-driven traffic flow models have recently emerged [87, 88]. Since these models are independent of any specific mathematical framework and incorporate randomness, using them to test a strategy can lead to a promising validation [54, 57].
III-F Scenarios for Strategy Validation
Figure 10: JAD performance in a more challenging multiple-wave scenario (sourced from [54]). Both (a) and (b) share the same inflow. Fewer waves are generated in (a) by setting a lower slowdown probably at the bottleneck, while more in (b).
As shown in Figure 1, it is quite common to observe that multiple and periodic stop-and-go waves appear on the freeway at the same time, and therefore, it is necessary to evaluate the strategies in a more practical and common scenario. Unfortunately, as shown in Table III, many strategies are validated in single-wave scenarios, which are relatively simple compared to the real-world observations. Single-wave scenarios are well-suited for initially testing the strategies, while multiple-wave scenarios allow for a more thorough assessment of the strategies in a setting closer to real-world conditions. Obviously, multiple-wave scenarios present greater challenges (Figure 10) [54, 56]. For the cell-based MPC strategies, the model resolution, i.e., the cell size, may not be sufficiently fine to distinguish multiple waves. However, it remains uncertain whether the current cell size is sufficient to capture multiple waves, or what level of granularity would be appropriate [62]. Even for the SPECIALIST, although it was designed for the single (or isolated) wave scenario, it still deserves a test in a more realistic scenario such as the one shown in Figure 1(d) [45]. Then, new problems can be uncovered, inspiring new research questions and corresponding solutions.
@@ III-F Scenarios for Strategy Validation
III-G Field Test and Practical Deployment
III-G1 Variable Speed Limit
@@
Figure 11: Two ways to disseminate speed limits (The background image is sourced from [21]): (a) One-shot plan generated by SPECIALIST, where the VSL displayed across multiple VMS is scheduled in advance based on the time-space plan shown in the figure. Note that controlling the yellow region is not theoretically required but is practically preferred, particularly to extend the speed limits into area 2 downstream of line CD. It would feel illogical to drivers if the speed limit is lifted just as they enter the jam. (b) bi-VMS configuration.
At present, the VSL strategies require the installation of a series of gantries and VMS for support. Referring to JAD strategies, it may be feasible to propose a low-cost bi-VSL strategy, in which the upstream VSL reduces the speed limit, while the downstream one is responsible for setting the speed back to normal (Figure 11(b)). The bi-location VSL might not only be more practical but also sufficient to set a low-speed region to prevent stop-and-go wave propagation. Still, it is worth noting that—similar to JAD, as discussed at the end of Section II-B—the stability behind the blue line in Figure 11(b) remains unclear and warrants further investigation.
III-G2 Jam Absorption Driving
@@
First, the behavior similar to JAD has occurred in the real world. As shown in Figure 12, a police car was recorded swerving on the freeway in California, USA, to control the following traffic at a high speed. In Figure 12, it was 70 km/h, which is higher than the lowest VSL at 60 km/h required by many VSL studies. The real-world police car swerving highlights the feasibility of JAD. Although the goal of the police car in this example might not be jam absorption, it greatly inspires a potential real-world implementation of JAD. In addition, we are still not sure how often such swerving behavior happens in practice, but it is not difficult to find similar videos on the Internet.
@@
Regarding the traffic stability behind the JAD vehicle, unfortunately, no data is currently available for the traffic conditions behind the police car. It would be interesting to see whether the upstream flow remained stable, as this could offer important insights into the feasibility and effectiveness of single- or few-vehicle-based JAD.
@@
Second, empirical data suggest that maintaining an appropriate gap from the leading vehicle when joining a stop-and-go wave could be effective in suppressing it, as shown in Figure 13. Although one may argue that this observation is related to the nature drop in demand, it still implies the potential feasibility of JAD, in particular from the microscopic perspective. The key lies in identifying the applicable conditions under which JAD can be effectively implemented. Therefore, it suggests another promising research direction, which is to extract similar scenarios from the existing extensive empirical trajectory datasets and analyze them to identify the conditions under which such behavior could contribute to wave suppression, ultimately proposing an empirical observation-based JAD strategy or even other CAV-based strategies beyond JAD.
Figure 13: Empirical observations suggest that maintaining an appropriate gap from the leading vehicle when joining a stop-and-go wave may effectively suppress its propagation without triggering secondary waves. The regional screenshots were extracted from time-space diagrams of vehicle trajectories, which were plotted using data collected from the Hanshin Expressway Route 11 (Ikeda Route) in Osaka City, Japan [18].
@@ IV Concluding Remarks
Stop-and-go waves are the common pattern of congestion on freeways/highways worldwide. With their severe adverse effects on traffic efficiency, driving safety, and emissions, stop-and-go waves persistently affect human society on a broad time-space scale. Over decades of research and development, a variety of measures and strategies have emerged to suppress the propagation of stop-and-go waves, including mainly VSL and JAD strategies. Unfortunately, despite important similarities, such as the shared goals and underlying logic, a disconnection between the two areas is observed. On the occasion of the 20th anniversary of the first stop-and-go wave suppression strategy [39] being proposed, this review summarizes the progress of both VSL and JAD areas in detail, pinpoints the strengths and weaknesses of each (Figure 14), and seeks to bridge the two areas, identify research opportunities, and streamline research directions.
Figure 14: Authors’ subjective assessment of VSL vs. JAD development.
Learning from JAD strategies, we recommend that the VSL area to place more focus on thorough validation using more realistic (e.g., stochastic) microscopic traffic simulations and in more realistic scenarios. The secondary wave issue (particularly type-II and type-III), which has been widely reported in the JAD area and is highly likely to exist in the VSL area, urgently needs to be tested and addressed. To enhance traffic estimation and prediction, more advanced techniques, such as deep learning, could be incorporated. To increase the generalization of the strategies, it is beneficial to consider more diverse traffic patterns, as well as introduce the latest AI tools (e.g., LLMs) to reduce the workload of fine-tuning. To reduce the need for constructing continuous, high-cost gantries for VSL strategies, a lightweight bi-VSL strategy could be proposed, drawing inspiration from JAD strategies.
Compared to VSL strategies, the theoretical basis underlying the design of JAD strategies is less well developed. This may be the main reason why the secondary wave issue has been inadequately addressed, and some research may even be biased. The real-world test in [21] provides strong evidence for the effectiveness of VSL strategies. Unfortunately, no similar field test for JAD has been reported. Since they share similar logic, JAD should also work similarly to VSL. Although it is not directly related to stop-and-go wave suppression, the observation of a police car swerving on a freeway at high speeds provides us with great confidence that JAD is also practically feasible. Making the current requirements (i.e., model input) for JAD implementation more practical is a valuable research effort that we can make before a real-world test opportunity arises.
@@
Unlike other areas of transportation research that center on (e.g.) understanding mechanisms or identifying patterns, freeway traffic control directly interacts with the traffic system, placing exceptionally high demands on the robustness and adaptability of the models and strategies. Given the similarities in research questions and goals, it might be worth exploring a unified framework that decouples modeling from implementation. Specifically, at the theoretical level, we may focus on stop-and-go wave suppression-oriented traffic flow optimization, through, e.g., traffic-based optimal control with the goal of minimizing travel time or suppressing the waves. In implementation, choosing between VSL and JAD can be guided by their respective strengths and limitations, depending on the practical conditions, with the aim of achieving the most effective outcome. Such a plug-in modeling-implementation structure enhances the overall framework’s generality and flexibility.